Solve Me!


1. Manuel has a container that contains a sample of nitrogen gas and a tightly fitting movable piston that does not allow any of the gas to escape. During a thermodynamics process, 200 Joules of heat enter the gas, and the gas does 300 Joules of work in the process. What was the change in internal energy of the gas during the process described above?



2. The work done when a gas is compressed in a cylinder is 462 J. During this process, there is a heat transfer of 128 J from the gas to the surroundings. Calculate the energy change for this process.


1
Expert's answer
2021-05-17T04:11:10-0400

1.) The first law of thermodynamics states that the change in internal energy of a system equals the net heat transfer into the system minus the net work done by the system. In equation form, the first law of thermodynamics is ΔU = Q − W. Here ΔU is the change in internal energy U of the system.

2.) "\Delta E=q+w

 = -128+ 462

 = 334J"